# Imitative Membership Inference Attack

### v1

@@ 4.4 Ablation Study
temperature TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Impact of Imitative Training. We evaluate the effectiveness of imitative training by comparing $\mathsf{IMIA}$ with two variants: (i) $\mathsf{IMIA}$$_{\text{distill}}$, which replaces imitative training with standard model distillation using KL divergence as the loss function [20]; and (ii) $\mathsf{IMIA}$$_{\text{shadow}}$, which substitutes imitative training with target-agnostic shadow training. Table 7 reports results on MNIST and CIFAR-10. Both variants show substantial degradation across all metrics, especially in the low false-positive regime. Using model distillation reduces average performance by over 20%, while replacing our approach with shadow training cuts attack performance by at least half. These results demonstrate that effective imitation should capture the target’s membership signals, rather than its general behaviors across models.
@@
Selection $k$ TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Strategy TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Signal TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Impact of Different Membership Signals. We analyze how the choice of membership signal affects the performance of $\mathsf{IMIA}$. The signal is used both to train imitative models and to compute final scores for inference. We compare three options: (i) loss, which uses the output loss of a query instance; (ii) probability, our default approach, which uses the softmax output probabilities; (iii) pre-softmax, which uses the outputs of the final layer before softmax. While all black-box MIAs assume access only to probabilities, using pre-softmax activation is also common in the implementation of many attacks [3, 54, 60, 9]. Results for both adaptive and non-adaptive settings are shown in Table˜12. As observed, using the prediction loss significantly degrades performance, and using pre-softmax activation instead of probabilities consistently yields better results. We also find that using pre-softmax activation to compute scaled confidence scores is numerically more stable and computationally simpler, as it avoids the logarithmic operation on probability values. These results are consistent with previous studies [3], and we recommend using pre-softmax activation for our attack when available.
Impact of Model Architectures and Distribution Shift. We also examine the robustness of $\mathsf{IMIA}$ in two scenarios: (i) when the adversary is unaware of the target model’s architecture, and (ii) when there is a distribution mismatch between the adversary’s data and the target’s data. Experimental results show that mismatched model architectures have a minor impact on $\mathsf{IMIA}$’s performance. Distribution shifts would cause a noticeable performance drop for our attack, a trend also observed in prior work [3, 9, 60]. Due to the page limit, these experiments are detailed in Section˜B.2.
@@ 4.5 Additional Investigations
Method TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Method TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Attack on Non-Image Datasets. We evaluate $\mathsf{IMIA}$ on two widely-used non-image datasets, Purchase and Texas [48], using a multilayer perceptron (MLP) as the target model and comparing it with leading MIAs. Results in both non-adaptive and adaptive settings are presented in Table˜13 and Table˜14, respectively. As shown, $\mathsf{IMIA}$ consistently outperforms existing attacks, particularly in terms of TPR at low FPR. We notice that all evaluated MIAs perform worse on these non-image datasets than on image datasets like CIFAR-100, despite both having 100 classes. We attribute this discrepancy to the smaller generalization gap in the models trained on non-image data. For instance, as shown in Table˜18, our Purchase model shows a modest gap between its training and validation accuracies (97.7% vs. 79.5%), whereas the CIFAR-100 model exhibits a much larger gap indicative of significant overfitting (99.4% vs. 68.9%). Since MIAs typically exploit overfitting for attack, they are less effective against these better-generalized models. Similar observations have been found in prior work [3, 9].
@@
Table 16: Performance of adaptive attacks against a ResNet50 model trained on CIFAR-10 using DP-SGD.
@@
Attack Against DP-SGD. Differentially Private Stochastic Gradient Descent (DP-SGD) [1] is an effective defense mechanism against privacy attacks, including MIAs. Following prior work[3, 19], we evaluate the effectiveness of DP-SGD in defending against $\mathsf{IMIA}$. Specifically, we fix the clipping norm $C$ to 10 and test $\mathsf{IMIA}$ on a ResNet50 model trained on CIFAR-10 under the non-adaptive setting. As shown in Table˜15, even applying only gradient clipping (without noise addition) significantly reduces both model accuracy and the effectiveness of our attack. We also assess the impact of DP-SGD in the adaptive setting in Table˜16, and observe consistent trends: as the privacy guarantee strengthens, all attacks experience reduced performance, and their performance gaps narrow. Nevertheless, $\mathsf{IMIA}$ consistently outperforms existing MIAs, especially in the low false-positive regime.
@@ 5 Related Work
Neural networks have been known to be vulnerable to leaking sensitive information about their training datasets. Various attacks [4, 14, 16] have been proposed to quantify the extent of this data leakage and assess the associated risks. In this paper, we focus on the membership inference attack (MIA) [48], which aims to predict whether a specific data instance was included in a target model’s training set. MIA has become a widely used tool to audit the data privacy of machine learning models [39, 51, 41, 23, 10]. The first MIA against ML models was introduced by [48], who also proposed the technique of shadow training. MIAs and the shadow training have since been extended to various scenarios, including white-box [29, 40, 45], black-box [36, 35, 50, 46, 24, 31], label-only access [33, 6], and knowledge distillation [22] settings. Existing black-box MIAs can be categorized into the following two groups.
@@
Non-Adaptive MIAs. In this setting, the adversary trains shadow models (or avoids model training entirely) before accessing the membership queries. Several techniques have been introduced, such as score functions [59, 50], difficulty calibration [45, 53, 19], loss trajectory [30, 34], model distillation [15, 58], quantile regression [2], and hypothesis testing [60, 9]. However, these attacks typically rely on target-agnostic shadow training and require training hundreds of shadow models to achieve good performance.
@@
Adaptive MIAs. In the adaptive setting, the adversary can perform instance-by-instance behavioral analysis by training additional predictors [40] or conducting parametric hypothesis testing [3, 54]. A recent work [9] goes beyond individual instance attacks by exploring the membership dependence among query instances. However, these attacks suffer from significant computational costs, and recent studies [60, 2, 43] argue that such adaptive attacks are impractical, as they require training new models for each batch of queries.
@@
Model Distillation in MIAs. Model distillation [20, 25] has been used in MIAs in two main approaches: (i) it is used to train more reliable shadow out models for inference, as explored in [58, 15]; (ii) it serves as a tool to simulate the intermediate learning dynamics of the target model, and derive membership signals from this simulated learning process [30, 34]. However, these methods use standard distillation techniques and only model the target’s behavior on non-members. In contrast, we propose a novel imitative training approach tailored for MIAs, which mimics the target model’s behavior for both members and non-members, significantly enhancing both attack efficacy and efficiency.
@@
Evaluation of MIAs. Early studies [48, 59, 46] evaluated membership inference as a binary classification task using metrics like balanced accuracy. More recent research [3, 37] has emphasized evaluating attacks by true positive rate (TPR) at a very low false positive rate (FPR), and most MIAs perform poorly under this metric. In line with the current consensus in the MIA literature, we also adopt TPR at low FPRs as the primary evaluation metric in this paper.
@@ 6 Conclusion
In this paper, we introduce $\mathsf{IMIA}$, a new MIA built upon a novel imitative training technique to train target-informed imitative models for inference. Imitative training addresses the high variation of shadow models by strategically distilling membership signals from the target model. Extensive experiments in various attack settings demonstrate the effectiveness and efficiency of $\mathsf{IMIA}$. One limitation of our paper is that it mainly focuses on classification models, leaving the generalization to generative models [17, 11, 12] unexplored. Nonetheless, our work provides an efficient and practical attack, opening new directions for future research in MIAs.
@@ A.1 Dataset Description
MNIST. MNIST [28] is a benchmark dataset for evaluating handwritten digit classification algorithms. It contains 60,000 grayscale images of 28×28 pixels for training and 10,000 images for testing. The dataset consists of 10 classes, representing digits from 0 to 9.
Fashion-MNIST. Fashion-MNIST (FMNIST) [55] is a dataset designed as a more challenging replacement for MNIST, containing 60,000 grayscale images of size 28×28 pixels for training and 10,000 images for testing. It consists of 10 classes representing different fashion items, such as t-shirts, trousers, and sneakers.
CIFAR10. CIFAR-10 [26] is a benchmark dataset for general image classification tasks, containing 60,000 color images of size 32×32 pixels, equally distributed across 10 classes, including animals like cats, dogs, and birds, as well as vehicles like airplanes and trucks.
CIFAR-100. CIFAR-100 [26] is a dataset similar to CIFAR-10 but with a greater level of complexity, as it contains 100 classes instead of 10. It also includes 60,000 color images of size 32×32 pixels, with each class containing 600 images.
Purchase. Purchase [48] is a dataset of shopping records with 197,324 samples of 600 dimensions. Following previous works [48, 30, 3], we cluster these data into 100 classes to train a Multi-Layer Perceptron (MLP) classifier.
Texas. This dataset is to predict a patient’s primary medical procedure using 6,170 binary features. Following [48], the data comprises records from 67,330 patients, with the 100 most frequent procedures used as classification labels.
Table 17: Prediction accuracy of image datasets.
Model MNIST FMNIST CIFAR-10 CIFAR-100
Train Val Train Val Train Val Train Val
ResNet50 100.0% 99.5% 100.0% 94.8% 99.7% 90.4% 99.4% 68.9%
VGG16 100.0% 99.7% 100.0% 95.6% 99.9% 91.7% 99.6% 72.3%
DenseNet121 100.0% 99.6% 100.0% 96.1% 99.9% 90.2% 99.9% 71.5%
MobileNetV2 100.0% 99.5% 100.0% 94.9% 99.9% 97.1% 100.0% 72.4%
Table 18: Prediction accuracy of non-image datasets.
Model Purchase Texas
Train Val Train Val
MLP 97.7% 79.5% 96.5% 78.7%
Table 19: Data partitioning scheme. The query dataset $D_{\text{query}}$ and the auxiliary dataset $D_{\text{auxiliary}}$ are strictly disjoint. The target model is trained on a randomly sampled subset from $D_{\text{query}}$. The reference dataset is used by some MIAs [19, 30] to train their attack models.
@@ A.2 Accuracy of the Trained Target Model
We report the training and validation accuracies of the target model for both image and non-image datasets in Table˜17 and Table˜18, respectively.
@@ A.3 Number of Shadow Models of Baselines
A critical hyperparameter for shadow-based MIAs is the number of trained shadow models. For each baseline method, we adopted the number of shadow models recommended in the original paper, if available. When a number was not specified, we select the maximum number of shadow models until the attack performance saturates to ensure its optimal attack performance. Specifically, we use the following configurations: 10 shadow models for Calibration [53] and SeqMIA [30]; 128 shadow models for Canary [54], GLiRA [15], RMIA [60], RAPID [19]; and 256 shadow models for Attack-R [58], Attack-D [58], LiRA [3], and PMIA [9]. The number of shadow models is consistent across all datasets.

### v2

@@ 4.4 Ablation Study
temperature TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Impact of Imitative Training. We evaluate the effectiveness of imitative training by comparing $\mathsf{IMIA}$ with two variants: (i) $\mathsf{IMIA}$$_{\text{distill}}$, which replaces imitative training with standard model distillation using KL divergence as the loss function [arxiv15distill]; and (ii) $\mathsf{IMIA}$$_{\text{shadow}}$, which substitutes imitative training with target-agnostic shadow training. Table 7 reports results on MNIST and CIFAR-10. Both variants show substantial degradation across all metrics, especially in the low false-positive regime. Using model distillation reduces average performance by over 20%, while replacing our approach with shadow training cuts attack performance by at least half. These results demonstrate that effective imitation should capture the target’s membership signals, rather than its general behaviors across models.
@@
Selection $k$ TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Strategy TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Signal TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Impact of Different Membership Signals. We analyze how the choice of membership signal affects the performance of $\mathsf{IMIA}$. The signal is used both to train imitative models and to compute final scores for inference. We compare three options: (i) loss, which uses the output loss of a query instance; (ii) probability, our default approach, which uses the softmax output probabilities; (iii) pre-softmax, which uses the outputs of the final layer before softmax. While all black-box MIAs assume access only to probabilities, using pre-softmax activation is also common in the implementation of many attacks [sp22lira, iclr23canary, icml24rmia, ndss26cpmia]. Results for both adaptive and non-adaptive settings are shown in Table˜12. As observed, using the prediction loss significantly degrades performance, and using pre-softmax activation instead of probabilities consistently yields better results. We also find that using pre-softmax activation to compute scaled confidence scores is numerically more stable and computationally simpler, as it avoids the logarithmic operation on probability values. These results are consistent with previous studies [sp22lira], and we recommend using pre-softmax activation for our attack when available.
Additional Ablation Studies. We also evaluate the robustness of $\mathsf{IMIA}$ under varying query budgets, scenarios where the adversary is unaware of the target model’s architecture, and cases involving distribution mismatches between the adversary’s data and the target’s data. Due to space limitations, the results of these experiments are provided in Appendix B.2.
@@ 4.5 Additional Investigations
Method TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Method TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
Attack on Non-Image Datasets. We evaluate $\mathsf{IMIA}$ on two widely-used non-image datasets, Purchase and Texas [sp17miashokri], using a multilayer perceptron (MLP) as the target model and comparing it with leading MIAs. Results in both non-adaptive and adaptive settings are presented in Table˜13 and Table˜14, respectively. As shown, $\mathsf{IMIA}$ consistently outperforms existing attacks, particularly in terms of TPR at low FPR. We notice that all evaluated MIAs perform worse on these non-image datasets than on image datasets like CIFAR-100, despite both having 100 classes. We attribute this discrepancy to the smaller generalization gap in the models trained on non-image data. For instance, our Purchase model shows a modest gap between its training and validation accuracies (97.7% vs. 79.5%), whereas the CIFAR-100 model exhibits a much larger gap indicative of significant overfitting (99.4% vs. 68.9%). Since MIAs typically exploit overfitting for attack, they are less effective against these better-generalized models. Similar observations have been found in prior work [sp22lira, ndss26cpmia].
@@
Table 16: Performance of adaptive attacks against a ResNet model trained on CIFAR-10 using DP-SGD.
@@
Attack Against DP-SGD. Differentially Private Stochastic Gradient Descent (DP-SGD) [ccs16dpsgd] is an effective defense mechanism against privacy attacks, including MIAs. Following prior work[sp22lira, ccs24rapid], we evaluate the effectiveness of DP-SGD in defending against $\mathsf{IMIA}$. Specifically, we fix the clipping norm $C$ to 10 and test $\mathsf{IMIA}$ on a ResNet model trained on CIFAR-10 under the non-adaptive setting. As shown in Table˜15, even applying only gradient clipping (without noise addition) significantly reduces both model accuracy and the effectiveness of our attack. We also assess the impact of DP-SGD in the adaptive setting in Table˜16, and observe consistent trends: as the privacy guarantee strengthens, all attacks experience reduced performance, and their performance gaps narrow. Nevertheless, $\mathsf{IMIA}$ consistently outperforms existing MIAs, especially in the low false-positive regime.
@@ 5 Related Work
Neural networks have been known to be vulnerable to leaking sensitive information about their training datasets. Various attacks [usenix21llm_extract, ccs15inversion, ccs18property] have been proposed to quantify the extent of this data leakage and assess the associated risks. In this paper, we focus on the membership inference attack (MIA) [sp17miashokri], which aims to predict whether a specific data instance was included in a target model’s training set. MIA has become a widely used tool to audit the data privacy of machine learning models [arxiv20privacy_meter, tensorflow_mia, sp21auditing, nips20auditing, ccs25synmeter]. The first MIA against ML models was introduced by [sp17miashokri], who also proposed the technique of shadow training. MIAs and the shadow training have since been extended to various scenarios, including white-box [usenix20stolen, sp19comprehensive, icml19white], black-box [long2018understanding, liu2022ml, usenix21systematic, ndss19ml-leak, arixv20revisiting, codaspy21mia], label-only access [ccs21label, icml21label], and knowledge distillation [nips23distill] settings. Existing black-box MIAs can be categorized into the following two groups.
@@
Non-Adaptive MIAs. In this setting, the adversary trains shadow models (or avoids model training entirely) before accessing the membership queries. Several techniques have been introduced, such as score functions [csf18privacy, usenix21systematic], difficulty calibration [icml19white, iclr22calibrate, ccs24rapid], loss trajectory [ccs24seqmia, ccs22trajectory], model distillation [tifs25glira, ccs22enhanced], quantile regression [nips23quantile], and hypothesis testing [icml24rmia, ndss26cpmia]. However, these attacks typically rely on target-agnostic shadow training and require training hundreds of shadow models to achieve good performance.
@@
Adaptive MIAs. In the adaptive setting, the adversary can perform instance-by-instance behavioral analysis by training additional predictors [sp19comprehensive] or conducting parametric hypothesis testing [sp22lira, iclr23canary]. A recent work [ndss26cpmia] goes beyond individual instance attacks by exploring the membership dependence among query instances. However, these attacks suffer from high computational costs, and recent studies [icml24rmia, nips23quantile, usenix25free] argue that such adaptive attacks are impractical, as they require training new models for each batch of queries.
@@
Model Distillation in MIAs. Model distillation [arxiv15distill, ijcai21distill] has been used in MIAs in two main approaches: (i) it is used to train more reliable shadow out models for inference, as explored in [ccs22enhanced, tifs25glira]; (ii) it serves as a tool to simulate the intermediate learning dynamics of the target model, and derive membership signals from this simulated learning process [ccs24seqmia, ccs22trajectory]. However, these methods use standard distillation techniques and only model the target’s behavior on non-members. In contrast, we propose an imitative training approach tailored for MIAs, which mimics the target model’s behavior for both members and non-members, significantly enhancing both attack efficacy and efficiency.
@@
Evaluation of MIAs. Early studies [sp17miashokri, csf18privacy, ndss19ml-leak] evaluated membership inference as a binary classification task using metrics like balanced accuracy. More recent research [sp22lira, eurosp20pragmatic] has emphasized evaluating attacks by true positive rate (TPR) at a very low false positive rate (FPR), and most MIAs perform poorly under this metric. In line with the current consensus in the MIA literature, we also adopt TPR at low FPRs as the primary evaluation metric in this paper.
@@ 6 Conclusion
In this paper, we introduce $\mathsf{IMIA}$, a new MIA built upon a novel imitative training technique to train target-informed imitative models for inference. Imitative training addresses the high variation of shadow models by strategically distilling membership signals from the target model. Extensive experiments in various attack settings demonstrate the effectiveness and efficiency of $\mathsf{IMIA}$. One limitation of our paper is that it mainly focuses on classification models, leaving the generalization to generative models [arxiv17logan, icml23diffusion, arxiv24llmmia] unexplored. Nonetheless, our work provides an efficient and practical attack, opening new directions for future research in MIAs.
Ethical Considerations
Our work introduces Imitative Membership Inference Attack ($\mathsf{IMIA}$), a more powerful and efficient approach to membership inference. We recognize our responsibility to carefully assess its ethical implications. Our analysis follows the stakeholder-based framework recommended by USENIX Security and is guided by the principles of The Menlo Report.
Stakeholder-Based Analysis. We identify several key stakeholders impacted by this research:
  • Machine Learning Practitioners. This is our primary audience. We provide ML practitioners with a more effective tool for privacy auditing, allowing them to identify vulnerabilities in their models and improve privacy protections.
  • Data Subjects. The publication of a more effective MIA could theoretically be used to infer personal information, leading to privacy violations. However, our experiments exclusively use public, non-sensitive datasets, and we do not target any individuals or proprietary models. Moreover, by publishing these attacks and their potential mitigations, as described below, our work helps strengthen the overall privacy and security of the deep learning ecosystem.
  • Adversaries. Our work could potentially be used by adversaries to exploit privacy vulnerabilities. In recognition of this, we discuss the mitigations below to minimize this risk.
Mitigations. We have taken the following concrete steps to mitigate potential harms:
  • Exclusive Use of Public, Non-Sensitive Data. Our research was conducted exclusively on public datasets (e.g., CIFAR-10), which do not contain Personally Identifiable Information (PII). We did not use any private or sensitive data, nor did we target specific individuals or proprietary models, ensuring no one’s privacy was violated during our research.
  • Validation of Existing Defenses. We evaluate that established defenses like DP-SGD remain an effective mitigation strategy against $\mathsf{IMIA}$. By demonstrating that practical defenses exist, we lower the privacy risk introduced by our work and provide clear guidance for developers.
Justification for Research and Publication. The vulnerabilities that $\mathsf{IMIA}$ exploits are fundamental to current deep learning practices. It is highly likely that adversaries could independently discover similar techniques, and we believe it is crucial to proactively disclose this attack to the research community. By publishing our findings, we enable defenders to address the vulnerabilities in their models before they are exploited in real-world scenarios.
We emphasize that the responsible use of such attack methods should be used responsibly to strengthen the security and privacy of ML systems, rather than maliciously exploiting them. We encourage the research community to use our findings in a manner that promotes ethical research and enhances privacy protections for users and data subjects.
Open Science
Our artifact is a code repository containing the source code for our proposed attack, along with scripts required to replicate the experiments presented in this paper. A detailed README.md file within the repository provides step-by-step instructions for setting up the environment and running the experiments. Artifacts are available at https://github.com/zealscott/IMIA.
@@ A.1 Dataset Description
MNIST. MNIST [mnist] is a benchmark dataset for evaluating handwritten digit classification algorithms. It contains 60,000 grayscale images of 28×28 pixels for training and 10,000 images for testing. The dataset consists of 10 classes, representing digits from 0 to 9.
Fashion-MNIST. Fashion-MNIST (FMNIST) [fmnist] is a dataset designed as a more challenging replacement for MNIST, containing 60,000 grayscale images of size 28×28 pixels for training and 10,000 images for testing. It consists of 10 classes representing different fashion items, such as t-shirts, trousers, and sneakers.
CIFAR10. CIFAR-10 [cifar10] is a benchmark dataset for general image classification tasks, containing 60,000 color images of size 32×32 pixels, equally distributed across 10 classes, including animals like cats, dogs, and birds, as well as vehicles like airplanes and trucks.
CIFAR-100. CIFAR-100 [cifar10] is a dataset similar to CIFAR-10 but with a greater level of complexity, as it contains 100 classes instead of 10. It also includes 60,000 color images of size 32×32 pixels, with each class containing 600 images.
Purchase. Purchase [sp17miashokri] is a dataset of shopping records with 197,324 samples of 600 dimensions. Following previous works [sp17miashokri, ccs24seqmia, sp22lira], we cluster these data into 100 classes to train a Multi-Layer Perceptron (MLP) classifier.
Texas. This dataset is to predict a patient’s primary medical procedure using 6,170 binary features. Following [sp17miashokri], the data comprises records from 67,330 patients, with the 100 most frequent procedures used as classification labels.
Table 17: Prediction accuracy of used datasets.
Model MNIST FMNIST CIFAR-10 CIFAR-100 Purchase Texas
Train Val Train Val Train Val Train Val Train Val Train Val
ResNet 100.0% 99.5% 100.0% 94.8% 99.7% 90.4% 99.4% 68.9% - - - -
VGG16 100.0% 99.7% 100.0% 95.6% 99.9% 91.7% 99.6% 72.3% - - - -
DenseNet121 100.0% 99.6% 100.0% 96.1% 99.9% 90.2% 99.9% 71.5% - - - -
MobileNetV2 100.0% 99.5% 100.0% 94.9% 99.9% 97.1% 100.0% 72.4% - - - -
MLP - - - - - - - - 97.7% 79.5% 96.5% 78.7%
Table 18: Data partitioning scheme. The query dataset $D_{\text{query}}$ and the auxiliary dataset $D_{\text{auxiliary}}$ are strictly disjoint. The target model is trained on a randomly sampled subset from $D_{\text{query}}$. The reference dataset is used by some MIAs [ccs24rapid, ccs24seqmia] to train their attack models.
@@ A.2 Accuracy of the Trained Target Model
We use validation accuracy thresholds to determine whether a trained model needs to be rerun. Specifically, we require a minimum validation accuracy of 0.8 for MNIST and FMNIST, 0.6 for CIFAR-10, and 0.4 for CIFAR-100. If a model fails to meet the corresponding threshold, we retrain it using a different random seed. We report the training and validation accuracies of the target model in Table˜17.
@@ A.3 Number of Shadow Models of Baselines
A critical hyperparameter for shadow-based MIAs is the number of trained shadow models. For each baseline method, we adopted the number of shadow models recommended in the original paper, if available. When a number was not specified, we select the maximum number of shadow models until the attack performance saturates to ensure its optimal attack performance. Specifically, we use the following configurations: 10 shadow models for Calibration [iclr22calibrate] and SeqMIA [ccs24seqmia]; 128 shadow models for Canary [iclr23canary], GLiRA [tifs25glira], RMIA [icml24rmia], RAPID [ccs24rapid]; and 256 shadow models for Attack-R [ccs22enhanced], Attack-D [ccs22enhanced], LiRA [sp22lira], and PMIA [ndss26cpmia]. The number of shadow models is consistent across all datasets.